A car manufacturer claims that its car can accelerate from rest to 28 m/s in 14 seconds. Find the average acceleration and the distance the car has traveled while in the process of stopping


1
Expert's answer
2021-09-23T03:05:32-0400

Average Acceleration

"a=\frac{V_f-V_i}{t}"


"=(28-0)\/14=2m\/s^2"


distance=?


"d=st"

"d=28m/s \times14s"

"d=392m"